# Bayesian Inference General Procedures for A Single-subject Test Study

### v2

@@ 3 Simulation
3.1 Jefferys Prior Comparison Study
In this section, we conduct a comparative analysis of our proposed prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ against the priors $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$. We evaluate these priors under six different parameter settings for $(\alpha,\nu)$: $(-1,1)$, $(-10,10)$, $(-30,30)$, $(-50,50)$, $(-1,50)$, and $(-50,1)$. For each setting, the location and scale parameters are fixed at $\xi=-2,\omega=\sqrt{2}$. To simplify the notation, we use $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ as shorthand for $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$, $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$, and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$, respectively. For each parameter setting, we generate $N=1000$ samples, each with a sample size of $n=500$, from the non-linear skewed Student’s $t$ density function. We then numerically minimise the negative logarithm of the posterior distribution to obtain the maximum posterior estimators of $\alpha,\nu,\xi,\omega$. The performance of different priors is evaluated using the mean absolute deviation (MAD) from the true parameter. The MAD is calculated as $1/N\sum_{i=1}^{N}{\lvert\hat{\theta}_{i}-\theta\rvert}$, where $\theta$ is the true parameter and $\hat{\theta}_{i}$ is the $i$-th maximum posterior estimator of $\theta$. Our comparative study also includes the maximum likelihood estimators (MLE). The results are presented in Table 2.
As demonstrated in Table 2, when $|\alpha|$ and $\nu$ are less than 30, the performance of $\boldsymbol{\pi}^{J}$ aligns closely with that of $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$ in terms of estimating $\alpha$ and $\nu$. However, for larger values of $|\alpha|$ and $\nu$, $\boldsymbol{\pi}^{J}$ outperforms both $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$. Furthermore, $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ all outperform the maximum likelihood estimators (MLE) for larger $|\alpha|$ and $\nu$. In terms of estimating the location $\xi$ and scale $\omega$, the mean absolute deviations (MADs) of all priors are comparable to the MLE. These results underscore the efficacy of $\boldsymbol{\pi}^{J}$ for the non-linear skewed Student’s $t$ density function, particularly when $|\alpha|$ and $\nu$ exceed 30.
@@ 3.1 Jefferys Prior Comparison Study
Table 2: Comparison of the mean absolute deviation (MAD) among $\boldsymbol{\pi}^{J}$ with $c_{\nu}=0$ and $c_{\nu}=1$, $\boldsymbol{\pi}^{B}$, $\boldsymbol{\pi}^{D}$, and the Uniform prior. The maximum a posteriori (MAP) estimation with a Uniform prior distribution is generally equivalent to the maximum likelihood estimation (MLE). Each entry in the table represents the MAD, calculated based on 1000 replications.
@@ 3.2 Comparison with Existing Approaches
To evaluate the performance of BIGPAST in comparison to the $z$-score, $t$-score (Crawford and Howell, 1998), Crawford-Garthwaite Bayesian approach (Crawford and Garthwaite, 2007), and Anderson-Darling non-parametric approach (Scholz and Stephens, 1987) for abnormality detection in a single-subject against a control group, we adopt the “severe skew ($\gamma_{1}=-0.7$)” settings from Section 3.1.1 in Crawford et al. (2006). The parameter $\gamma_{1}$ is a measure of skewness in Crawford et al. (2006). This setting is equivalent to parameter settings: $\alpha=-3.23$, $\nu=7$, $\xi=0$, and $\omega=1$ in the definition of our skewed Student’s $t$ distribution. The single-subject and control group observations are generated from $f(x|-3.23,7,0,1)$. We consider sample sizes of 50, 100, 200, and 400 for the control group. The code for the simulations in this section can be found in GitHub repository: BIGPAST. Given a significance level of 0.05 and a one-sided test (less), we record the abnormality detection results using the $z$-score, $t$-score, Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST. This procedure is repeated one million times as Crawford et al. (2006). The results, presented in Table 3, indicate that BIGPAST outperforms the other methods in terms of false positive rate (FPR) and accuracy (ACC) when all single-subject observations are from the ground true distribution.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.070022 0.066415 0.066427 0.073271 0.051883
100 0.069632 0.067787 0.067786 0.056572 0.053742
200 0.066616 0.065660 0.065668 0.057564 0.052586
400 0.067888 0.067435 0.067436 0.056244 0.052051
ACC 50 0.929978 0.933585 0.933573 0.926729 0.948117
100 0.930368 0.932213 0.932214 0.943428 0.946258
200 0.933384 0.934340 0.934332 0.942436 0.947414
400 0.932112 0.932565 0.932564 0.943756 0.947949
@@
Table 3: The table presents the simulation results for false positive rate (FPR) and accuracy (ACC), derived from single-subject observations that are exclusively negative, with one-sided test (less). Each entry in the table is computed based on one million test outcomes. The abbreviations ‘CG’ and ‘AD’ denote the Crawford-Garthwaite Bayesian and Anderson-Darling methods, respectively.
@@
Given that all single-subject observations are from ground true distribution, implying a negative underlying result, the true positive rate and false negative rate (Type II error) are undefined. To address this, we modify the generation of single-subject observations such that 50% yield negative results and the remaining 50% yield positive results. These observations are generated using the method described in Section 3.4. The simulation is then conducted as previously described. The true positive rate and false negative rate for the $z$ score, $t$ score, Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST, pertaining to a two-sided test, are presented in Table 4. Similar simulations have been conducted for one-sided test (greater) and one-sided test (less), with the results detailed in Appendix, see Tables 7 and 8.
@@
From Table 4, we can conclude that BIGPAST outperforms the $t$-score and Crawford-Garthwaite methods in terms of false positive rate (FPR) and accuracy (ACC) when the single-subject observations are mixed. As normal distribution is a special case of skewed Student’s $t$ distribution ($\alpha=0,\nu=+\infty$), the performance of BIGPAST is still as good as those of $t$-score and Crawford-Garthwaite methods when the underlying distribution is normal distribution, see Table 9 in Appendix. The results are not surprising because the errors in this context originate from two primary sources: model misspecification and inherent randomness. Notably, BIGPAST is capable of effectively mitigating the errors arising from model misspecification; see Figure 3 in the next section.
$n$ $z$ $t$ CG AD BIGPAST
FPR 50 0.1320 0.1182 0.1182 0.1107 0.0428
100 0.1269 0.1199 0.1199 0.0782 0.0349
200 0.1207 0.1166 0.1167 0.0644 0.0334
400 0.1236 0.1215 0.1215 0.0368 0.0187
TPR 50 0.6908 0.6502 0.6502 0.9101 0.8327
100 0.6597 0.6415 0.6416 0.9356 0.8992
200 0.6420 0.6312 0.6311 0.9655 0.9552
400 0.6316 0.6264 0.6264 0.9660 0.9413
ACC 50 0.7794 0.7660 0.7660 0.8997 0.8949
100 0.7664 0.7608 0.7609 0.9287 0.9322
200 0.7606 0.7573 0.7572 0.9505 0.9609
400 0.7540 0.7525 0.7524 0.9646 0.9613
@@
Table 4: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under two-sided test. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’ and ‘AD’ represent the Crawford-Garthwaite Bayesian and the Anderson-Darling methods, respectively.
@@ 3.3 Model Misspecification Error
In this section, we conduct a comparative analysis of the BIGPAST, CG, and AD methodologies under varying alternative hypotheses and parameter configurations. The predetermined significance level is 0.05. The skewness parameter $\alpha$ and degree of freedom $\nu$ are assigned values from the sets $(1,10)$, $(1,5)$, $(2,5)$, and $(3,5)$, respectively. The location and scale of skewed Student’s $t$ distribution are set to 0 and 1, respectively. Other sample settings include a sample size $n=100$, $c_{\nu}=1$, and a Metropolis-Hastings sampling size of 2000. The comparison procedures are as follows: Firstly, $N=200$ independent samples are randomly generated from the skewed Student’s $t$ distribution with parameters $\alpha$ and $\nu$. Secondly, $m=1000$ single-subjects are randomly drawn, and the BIGPAST, CG, and AD tests are conducted on each of the $N$ samples. Thirdly, for each of $N$ samples, the Type I error (false positive rate), power (true positive rate), and accuracy are calculated based on the $m=1000$ single-subjects. Finally, the Type I error, power, and accuracy are reported in the form of boxplots for the four skewed Student’s $t$ parameter settings, as shown in Figures 10, 11, 12 and 2.
Under the mechanism of these parameter settings, the model is misspecified for the CG test. The theoretical assumption of CG test is the normal distribution, the mean and variance can be directly computed from the skewed Student’s $t$ distribution by $\mu=\delta b_{\nu},\nu>1$ and $\sigma^{2}=\nu/(\nu-2)-(b_{\nu}\delta)^{2},\nu>2$, where
@@
We utilise the total variation distance to quantify the disparity between skewed Student’s $t$ (for BIGPAST) and normal (for CG) distributions. The total variation distances of the aforementioned four settings are: 0.057, 0.111, 0.155 and 0.186, respectively, as illustrated in Figures 10(a), 11(a), 12(a) and 2(a). In the context of a two-sided hypothesis test, the divergence between the BIGPAST and CG methodologies becomes larger as the total variation distance increases, as shown in Figures 10, 11, 12 and 2. The results of the AD approach demonstrate robustness for different total variation distances, but perform significantly worse than BIGPAST in terms of Type I error, as seen in Figure 2(b). The AD approach also performs slightly worse than BIGPAST in terms of accuracy, as depicted in Figure 2(d).
@@
Figure 2: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=3$ and $\nu=5$.  Figure 2(a) shows the densities of skewed Student’s $t$ and normal distributions, the $\mu$ and $\sigma^{2}$ of normal distribution are equal to the mean and variance of skewed Student’s $t$ distribution. In fact, the normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student’s $t$ distribution. AD test is a nonparametric test; there is one result that is compared in the ‘two-sided’ category. TV distance is short for total variation distance. Each boxplot summarises over 200 independent replications.
The complexity of the one-sided test (either ‘greater’ or ‘less’) compared to the two-sided test due to the skewed Student’s $t$ distribution increases due to skewed Student’s $t$ distribution has light and heavy tails. The model misspecification error impacts the tails of the skewed Student’s $t$ distribution differently. When $\alpha>0$ (resp. $\alpha<0$), the skewed Student’s $t$ distribution has a right (resp. left) heavy tail. In this paragraph, we explore the scenario where the direction of the alternative hypothesis is ‘greater’, i.e., one-sided test (greater). If we implement the CG test, the 95% quantile of normal distribution is less than the 95% quantile of skewed Student’s $t$ distribution. Consequently, the Type I error of the CG test exceeds that of the BIGPAST test, leading to a higher power for the CG test, as shown in Figures 2(b) and 2(c). This can be explained by the light-red region between the 95% quantile of skewed Student’s $t$ distribution and the 95% quantile of normal distribution, as depicted in Figure 3(b). Theoretically, all single-case observations from this region will be classified as positive by the CG test and as negative by the BIGPAST test, leading us to refer to this region as the false positive region. The error generated by the false positive region arises from model misspecification rather than sampling randomness, as illustrated in Figure 3.
Conversely, for the light tail of the skewed Student’s $t$ distribution, the 5% quantile of the normal distribution is less than the 5% quantile of the skewed Student’s $t$ distribution. Given that the alternative is ‘less’ and $\alpha>0$, the region between the 5% quantile of skewed Student’s $t$ distribution and the 5% quantile of the normal distribution is referred to as the false negative region other than the false positive region. Because all the single-case observations from this region will be classified as negative by the CG test, but as positive by the BIGPAST test. Consequently, the theoretical Type I error is zero, excluding the random error of sampling, as depicted in Figure 2(b). The power of the CG test, i.e., the actual positive rate, is lower than that of the BIGPAST test, as illustrated in Figure 2(c).
@@
For the left skewed Student’s $t$ distribution ($\alpha<0$), we can obtain similar results just by reflecting the densities in line $x=0$. This will not be detailed here. From the preceding discussion, we can conclude that (1) The difference between the BIGPAST and CG methodologies becomes more significant in terms of Type I error, power, and accuracy as the total variation distance increases; (2) Assuming a skewed Student’s $t$ distribution, the BIGPAST approach outperforms the CG approach in terms of accuracy due to the presence of model misspecification error; (3) Given the asymmetry between the light and heavy tails of the skewed Student’s $t$ distribution, we recommend using a two-sided test instead of a one-sided test. This is because the combinations of a one-sided test (‘less’ or ‘greater’) and the sign of $\alpha$ significantly influence the evaluation of detection.

### v4

@@ 3 Simulation
3.1 Jeffreys Prior Comparison Study
In this section, we conduct a comparative analysis of our proposed prior $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$ against the priors $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$ and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$. We evaluate these priors under six different parameter settings for $(\alpha,\nu)$: $(-1,1)$, $(-10,10)$, $(-30,30)$, $(-50,50)$, $(-1,50)$, and $(-50,1)$. For each setting, the location and scale parameters are fixed at $\xi=-2,\omega=\sqrt{2}$. To simplify the notation, we use $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ as shorthand for $\boldsymbol{\pi}^{J}(\alpha,\nu,\xi,\omega)$, $\boldsymbol{\pi}^{B}(\alpha,\nu,\xi,\omega)$, and $\boldsymbol{\pi}^{D}(\alpha,\nu,\xi,\omega)$, respectively. For each parameter setting, we generate $N=1000$ samples, each with a sample size of $n=500$, from the skewed Student $t$ density function. We then numerically minimise the negative logarithm of the posterior distribution to obtain the maximum posterior estimators of $\alpha,\nu,\xi,\omega$. The performance of different priors is evaluated using the mean absolute deviation (MAD) from the true parameter. Alternative metric distances may also be applicable in this context; however, a detailed discussion of these alternatives is beyond the scope of this paper. The MAD is calculated as $1/N\sum_{i=1}^{N}{\lvert\hat{\theta}_{i}-\theta\rvert}$, where $\theta$ is the true parameter and $\hat{\theta}_{i}$ is the $i$-th maximum posterior estimator of $\theta$. Our comparative study also includes the maximum likelihood estimators (MLE). The results are presented in Table 2.
As demonstrated in Table 2, when $|\alpha|$ and $\nu$ are less than 30, the performance of $\boldsymbol{\pi}^{J}$ aligns closely with that of $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$ in terms of estimating $\alpha$ and $\nu$. However, for larger values of $|\alpha|$ and $\nu$, $\boldsymbol{\pi}^{J}$ outperforms both $\boldsymbol{\pi}^{B}$ and $\boldsymbol{\pi}^{D}$. Furthermore, $\boldsymbol{\pi}^{J}$, $\boldsymbol{\pi}^{B}$, and $\boldsymbol{\pi}^{D}$ all outperform the maximum likelihood estimators (MLE) for larger $|\alpha|$ and $\nu$. In terms of estimating the location $\xi$ and scale $\omega$, the mean absolute deviations (MADs) of all priors are comparable to the MLE. These results underscore the efficacy of $\boldsymbol{\pi}^{J}$ for the skewed Student $t$ density function, particularly when $|\alpha|$ and $\nu$ exceed 30.
Table 2: Comparison of the mean absolute deviation (MAD) among $\boldsymbol{\pi}^{J}$ with $c_{\nu}=0$ and $c_{\nu}=1$, $\boldsymbol{\pi}^{B}$, $\boldsymbol{\pi}^{D}$, and the Uniform prior. The estimation with a Uniform prior distribution is generally equivalent to the maximum likelihood estimation (MLE). Each entry in the table represents the MAD, calculated based on 1000 replications.
@@ 3.2 Comparison with Existing Approaches
To evaluate the performance of BIGPAST in comparison to the $z$-score, $t$-score (Crawford and Howell, 1998), Crawford-Garthwaite Bayesian approach (Crawford and Garthwaite, 2007), and Anderson-Darling non-parametric approach (Scholz and Stephens, 1987) for abnormality detection in a single-subject against a control group, we adopt the “severe skew ($\gamma_{1}=-0.7$)” settings from Section 3.1.1 in Crawford et al. (2006). The parameter $\gamma_{1}$ is a measure of skewness in Crawford et al. (2006). This setting is equivalent to parameter settings: $\alpha=-3.23$, $\nu=7$, $\xi=0$, and $\omega=1$ in the definition of our skewed Student $t$ distribution. The single-subject and control group observations are generated from $f(x|-3.23,7,0,1)$. We consider sample sizes of 50, 100, 200, and 400 for the control group. The simulation code in this section can be found in GitHub repository: BIGPAST. Given a significance level of 0.05, a one-sided test (less) and $\mathbf{c:d=100:0}$, we record the abnormality detection results using the $z$-score, $t$-score, hyperbolic arcsine transformation ($\log(x+\sqrt{x^{2}+1}),x\in\mathbb{R}$), Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST. Other transformations for $x\in\mathbb{R}$ are applicable; however, determining the optimal transformation is beyond the scope of this paper. Therefore, we will only consider the hyperbolic arcsine transformation in this study. This procedure is repeated one million times as Crawford et al. (2006). The results, presented in Table 3, indicate that BIGPAST outperforms the other methods in terms of false positive rate (FPR) and accuracy (i.e., ACC=(TP+TN)/(TP+TN+FP+FN).) when all single-subject observations are from the ground truth distribution.
@@
Table 3: The table presents the simulation results for false positive rate (FPR) and accuracy (ACC), derived from single-subject observations that are exclusively negative, with a one-sided test (less). Each entry in the table is computed based on one million test outcomes. The abbreviations ‘CG’, ‘CG-HA’ and ‘AD’ denote Crawford-Garthwaite Bayesian method, Crawford-Garthwaite Bayesian method based on hyperbolic arcsine transformation, and Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG-HA CG AD BIGPAST
FPR 50 0.0700 0.0664 0.0544 0.0664 0.0733 0.0519
100 0.0696 0.0678 0.0550 0.0678 0.0566 0.0537
200 0.0666 0.0657 0.0567 0.0657 0.0576 0.0526
400 0.0679 0.0674 0.0554 0.0674 0.0562 0.0521
ACC 50 0.9300 0.9336 0.9456 0.9336 0.9267 0.9481
100 0.9304 0.9322 0.9450 0.9322 0.9434 0.9463
200 0.9334 0.9343 0.9433 0.9343 0.9424 0.9474
400 0.9321 0.9326 0.9446 0.9326 0.9438 0.9479
@@
Given that all single-subject observations are from ground truth distribution, implying a negative underlying result, the true positive rate and false negative rate (Type II error) are undefined. To address this, we modify the generation of single-subject observations such that 50% yield negative results and the remaining 50% yield positive results, i.e., $c:d=50:50$. Even in the case of $c:d=100:0$,  Table 3 demonstrates that BIGPAST outperforms the other methods. To balance the trade-off between the false positive rate (FPR) and the true positive rate (TPR), we will focus on the case of $c:d=50:50$ for the remainder of this section. Additionally, results for the case of $c:d=80:20$ are provided in Appendix, see Table 9.
The observations are generated using the method described in Section 3.4. The true positive rate and false positive rate for the $z$ score, $t$ score, hyperbolic arcsine transformation, Crawford-Garthwaite Bayesian method, Anderson-Darling method, and BIGPAST, pertaining to a two-sided test, are presented in Table 4. Similar simulations have been conducted for the one-sided test (greater) and one-sided test (less), with the results detailed in Appendix, see Tables 7 and 8.
@@
From Table 4, we can conclude that BIGPAST outperforms the $t$-score and Crawford-Garthwaite methods in terms of false positive rate (FPR) and accuracy (ACC) when the single-subject observations are mixed. As Normal distribution is a special case of skewed Student $t$ distribution ($\alpha=0,\nu\to+\infty$), the performance of BIGPAST is still as good as those of $t$-score and Crawford-Garthwaite methods when the underlying distribution is the Normal distribution, see Table 10 in Appendix. The results are not surprising because the errors in this context originate from two primary sources: model misspecification and inherent randomness. Notably, BIGPAST can effectively mitigate the errors arising from model misspecification; see Figure 4 in the next section.
@@
Table 4: The simulation results of false positive rate (FPR), true positive rate (TPR) and accuracy (ACC), derived from single-subject observations that consist of 50% positive and 50% negative, are presented under a two-sided test. Each cell in the resulting table is calculated based on one million test outcomes. ‘CG’, ‘CG-HA’ and ‘AD’ denote the Crawford-Garthwaite Bayesian method, the Crawford-Garthwaite Bayesian method based on hyperbolic arcsine transformation and the Anderson-Darling methods, respectively.
$n$ $z$ $t$ CG-HA CG AD BIGPAST
FPR 50 0.1320 0.1182 0.0450 0.1182 0.1107 0.0428
100 0.1269 0.1199 0.0569 0.1199 0.0782 0.0349
200 0.1207 0.1166 0.0333 0.1167 0.0644 0.0334
400 0.1236 0.1215 0.0338 0.1215 0.0368 0.0187
TPR 50 0.6908 0.6502 0.8595 0.6502 0.9101 0.8327
100 0.6597 0.6415 0.8917 0.6416 0.9356 0.8992
200 0.6420 0.6312 0.9196 0.6311 0.9655 0.9552
400 0.6316 0.6264 0.9229 0.6264 0.9660 0.9413
ACC 50 0.7794 0.7660 0.9072 0.7660 0.8997 0.8949
100 0.7664 0.7608 0.9174 0.7609 0.9287 0.9322
200 0.7606 0.7573 0.9432 0.7572 0.9505 0.9609
400 0.7540 0.7525 0.9445 0.7524 0.9646 0.9613
@@ 3.3 Model Misspecification Error
In this section, we conduct a comparative analysis of the BIGPAST, CG, and AD methodologies under varying alternative hypotheses and parameter configurations. The predetermined significance level is 0.05. The skewness parameter $\alpha$ and degrees of freedom $\nu$ are assigned values from the sets $(1,10)$, $(1,5)$, $(2,5)$, and $(3,5)$, respectively. The location and scale of the skewed Student $t$ distribution are set to 0 and 1, respectively. Other sample settings include a sample size $n=100$, $c_{\nu}=1$, and a Metropolis-Hastings sampling size of 2000. The comparison procedures are as follows: Firstly, $N=200$ independent samples are randomly generated from the skewed Student $t$ distribution with parameters $\alpha$ and $\nu$. Secondly, $m=1000$ single-subjects are randomly drawn, and the BIGPAST, CG, and AD tests are conducted on each of the $N$ samples. Thirdly, for each of $N$ samples, the Type I error (false positive rate), power (true positive rate), and accuracy are calculated based on the $m=1000$ single-subjects. Finally, the Type I error, power, and accuracy are reported in the form of boxplots for the four skewed Student $t$ parameter settings, as shown in Figures 11, 12, 13 and 2.
Under the mechanism of these parameter settings, the model is misspecified for the CG test. The theoretical assumption of CG test is the Normal distribution, the mean and variance can be directly computed from the skewed Student $t$ distribution by $\mu=\delta b_{\nu},\nu>1$ and $\sigma^{2}=\nu/(\nu-2)-(b_{\nu}\delta)^{2},\nu>2$, where
@@
We utilise the total variation distance to quantify the disparity between skewed Student $t$ (for BIGPAST) and Normal (for CG) distributions. The total variation distances of the aforementioned four settings are: 0.057, 0.111, 0.155 and 0.186, respectively, as illustrated in Figures 11(a), 12(a), 13(a) and 2(a). In the context of a two-sided hypothesis test, the divergence between the BIGPAST and CG methodologies becomes larger as the total variation distance increases, as shown in Figures 11, 12, 13 and 2. The result of the AD approach demonstrates robustness for different total variation distances but performs significantly worse than BIGPAST in terms of Type I error, as seen in Figure 2(b). The AD approach also performs slightly worse than BIGPAST in terms of accuracy, as depicted in Figure 2(d).
@@
Figure 2: The comparison results of BIGPAST, CG and AD approaches under three alternative hypotheses: ‘two-sided’, ‘less’ and ‘greater’ respectively when $\alpha=3$ and $\nu=5$Figure 2(a) shows the densities of skewed Student $t$ and Normal distributions, the $\mu$ and $\sigma^{2}$ of Normal distribution are equal to the mean and variance of skewed Student $t$ distribution. In fact, the Normal distribution is the theoretical assumption of the CG test when the sample comes from the skewed Student $t$ distribution. TV distance is short for total variation distance. Each boxplot summarises over 200 independent replications.
The complexity of the one-sided test (either ‘greater’ or ‘less’) compared to the two-sided test due to the skewed Student $t$ distribution increases due to skewed Student $t$ distribution having light and heavy tails. The model misspecification error impacts the tails of the skewed Student $t$ distribution differently. When $\alpha>0$ (resp. $\alpha<0$), the skewed Student $t$ distribution has a right (resp. left) heavy tail. In this paragraph, we explore the scenario where the direction of the alternative hypothesis is ‘greater’, i.e., one-sided test (greater). If we implement the CG test, the 95% quantile of Normal distribution is less than the 95% quantile of skewed Student $t$ distribution. Consequently, the Type I error of the CG test exceeds that of the BIGPAST test, leading to a higher power for the CG test, as shown in Figures 2(b) and 2(c). This can be explained by the light-red region between the 95% quantile of skewed Student $t$ distribution and the 95% quantile of Normal distribution, as depicted in Figure 4(b). Theoretically, all single-case observations from this region will be classified as positive by the CG test and as negative by the BIGPAST test, leading us to refer to this region as the false positive region. The error generated by the false positive region arises from model misspecification rather than sampling randomness, as illustrated in Figure 4.
Conversely, for the light tail of the skewed Student $t$ distribution, the 5% quantile of the Normal distribution is less than the 5% quantile of the skewed Student $t$ distribution. Given that the alternative is ‘less’ and $\alpha>0$, the region between the 5% quantile of skewed Student $t$ distribution and the 5% quantile of the Normal distribution is referred to as the false negative region other than the false positive region. Because all the single-case observations from this region will be classified as negative by the CG test but as positive by the BIGPAST test. Consequently, the theoretical Type I error is zero, excluding the random error of sampling, as depicted in Figure 2(b). The power of the CG test, i.e., the actual positive rate, is lower than that of the BIGPAST test, as illustrated in Figure 2(c).
@@
(a) Alternative: ‘less’
(b) Alternative: ‘greater’
For the left skewed Student $t$ distribution ($\alpha<0$), we can obtain similar results just by reflecting the densities in line $x=0$. This will not be detailed here. From the preceding discussion, we can conclude that (1) The difference between the BIGPAST and CG methodologies becomes more significant in terms of Type I error, power, and accuracy as the total variation distance increases; (2) Assuming a skewed Student $t$ distribution, the BIGPAST approach outperforms the CG approach in terms of accuracy due to the presence of model misspecification error; (3) Given the asymmetry between the light and heavy tails of the skewed Student $t$ distribution, we recommend using a two-sided test instead of a one-sided test. This is because the combinations of a one-sided test (‘less’ or ‘greater’) and the sign of $\alpha$ significantly influence the evaluation of detection.
For the left skewed Student $t$ distribution ($\alpha<0$), we can obtain similar results just by reflecting the densities in line $x=0$. This will not be detailed here. From the preceding discussion, we can conclude that (1) The difference between the BIGPAST and CG methodologies becomes more significant in terms of Type I error, power, and accuracy as the total variation distance increases; (2) Assuming a skewed Student $t$ distribution, the BIGPAST approach outperforms the CG approach in terms of accuracy due to the presence of model misspecification error; (3) Given the asymmetry between the light and heavy tails of the skewed Student $t$ distribution, we recommend using a two-sided test instead of a one-sided test. This is because the combinations of a one-sided test (‘less’ or ‘greater’) and the sign of $\alpha$ significantly influence the evaluation of detection.